OpenStudy (anonymous):

7x - 7 = 13 + 12x I got 4. Is that correct?

OpenStudy (anonymous):

7x-7 = 13 +12x -7 = 13 + 5x -20 = 5x -4 = x Almost. Don't forget your signs.

OpenStudy (anonymous):

rats!

OpenStudy (anonymous):

x=-4

OpenStudy (saifoo.khan):

lol.